a rectangle has a length of 7cm and a width of 5cm. the rectangle will be widened by a scale factor of 6. calculates the length of the enlargement
Answer by ikleyn(52781)

.

area of the enlarged rectangle = (7*6)*(5*6) = 35*36 = 1260 cm^2 = 


     = 36 times (i.e.. 6^2 times) the area of the original triangle.    ANSWER